OpenStudy (anonymous):

Solve and check. 2^3x = 15

OpenStudy (anonymous):

ln 2^3x = ln15 (3x) ln 2 = ln 15 3x= (ln 15)/( ln 2) x = 1.30

OpenStudy (anonymous):

\[3x=\frac{\ln(15)}{\ln(2)}\] by the change of base formula then \[x=\frac{\ln(15)}{3\ln(2)}\] then a calculator

Directrix (directrix):

Is 2^3x = 15 meant to be (2^3) x = 15 OR 2^ (3x) = 15?

OpenStudy (anonymous):

i get 1.3 rounded, not 1.5